Bibliography

book

23/03-04 16:19

[abelson1985a] Harold Abelson and Gerald Jay Sussman. Structure and Interpretation of Computer Programs. The MIT Electrical Engineering and Computer Science Series. The MIT Press/McGraw-Hill Book Company, 1985.

[acm1987a] ACM. ACM Turing Award Lectures — The First Twenty Years: 1866–1985. ACM Press Anthology Series. ACM Press, 1987.

[adobe1985a] Adobe Systems Incorporated. PostScript Language — Reference Manual. Addison-Wesley, 1985.

[ahmed2002a] Khawar Zaman Ahmed and Cary E. Umrysh. Developing Enterprise Java Applications with J2EE and UML. Object Technology Series. Addison-Wesley, 2002.

[aho1983a] Alfred V. Aho, John E. Hopcroft, and Jeffrey D. Ullman. Data Structures and Algorithms. Addison-Wesley Series in Computer Science and Information Processing. Addison-Wesley, 1983.

[aho1986a] Alfred V. Aho, Ravi Sethi, and Jeffrey D. Ullman. Compilers — Principles, Techniques, and Tools. Addison-Wesley, 1986.

[aho1992a] Alfred V. Aho and Jeffrey D. Ullman. Foundations of Computer Science. Principles of Computer Science Series. Computer Science Press, 1992.

[anderson2001a] Ross Anderson. Security Engineering. Wiley Computer Publishing. John Wiley & Sons, 2001.

[andreasson1993a] Torbjørn Andréasson and Jan Skansholm. Getting Started with UNIX and X. Addison-Wesley, 1993.

[andrews1991a] Gregory R. Andrews. Concurrent Programming — Principles and Practice. Benjamin/Cummings, 1991.

[anton1984a] Howard Anton. Elementary Linear Algebra. John Wiley & Sons, 4 edition, 1984.

[arnow1998a] David Arnow and Gerald Weiss. Introduction to Programming using Java: An Object-Oriented Approach. Addison-Wesley, 1998.

[ayers1999a] Danny Ayers, Hans Bergsten, Michael Bogovich, Jason Diamond, Matthew Ferris, Marc Fleury, Ari Halberstadt, Paul Houle, Piroz Mohseni, Andrew Patzer, Ron Phillips, Sing Li, Krishna Vedati, Mark Wilcox, and Stefan Zeiger. Professional Java Server Programming. Wrox Press Ltd, 1999.

[bach1986a] Maurice J. Bach. The Design of the UNIX Operating System. Prentice Hall, 1986.

[bacon1998a] Jean Bacon. Concurrent Systems. International Computer Science Series. Addison-Wesley, 2 edition, 1998.

[balen2000a] Henry Balen. Distributed Object Architectures with CORBA. Managing Object Technology Series. Cambridge University Press/SIGS books, 2000.

[barnes1989a] J. G. P. Barnes. Programming in Ada. International Computer Science Series. Addison-Wesley, 3 edition, 1989.

[barr1990a] Michael Barr and Charles Wells. Category Theory for Computing Science. Prentice Hall International Series in Computer Science. Prentice Hall, 1990.

[barr1995a] Michael Barr and Charles Wells. Category Theory for Computing Science. Prentice Hall International Series in Computer Science. Prentice Hall, 2 edition, 1990.

[barret2001a] Daniel J. Barret and Richard E. Silverman. SSH: The Secure Shell. A Nutshell Handbook. O’Reilly & Associates, Inc., 2001.

[beazley2000a] David M. Beazley. Python Essential Reference. New Riders, 2000.

[bechtolsheim1993a] Stephan von Bechtolsheim. TEX in Practice, volume I: Basics of Monographs in Visual Communication. Springer-Verlag, 1993.

[bechtolsheim1993b] Stephan von Bechtolsheim. TEX in Practice, volume II: Paragraphs, Math, and Fonts of Monographs in Visual Communication. Springer-Verlag, 1993.

[bechtolsheim1993c] Stephan von Bechtolsheim. TEX in Practice, volume III: Tokens, Macros of Monographs in Visual Communication. Springer-Verlag, 1993.

[bechtolsheim1993d] Stephan von Bechtolsheim. TEX in Practice, volume IV: Output Routines, Tables of Monographs in Visual Communication. Springer-Verlag, 1993.

[bell1987a] Dough Bell, Ian Morrey, and John Pugh. Software Engineering — A Programming Approach. Prentice Hall, 1978.

[ben-ari1990a] M. Ben-Ari. Principles of Concurrent and Distributed Programming. Prentice Hall International Series in Computer Science. Prentice Hall, 1990.

[bennet1987a] J. Michael Bennet. 68000 Assembly Language Programming. Prentice Hall, 1987.

[bentley1988a] John P. Bentley. Principles of Measurement Systems. Longman Scientific & Technical, 2 edition, 1988.

[bernstein1997a] Philip A. Bernstein. Principles of Transaction Processing. The Morgan Kaufmann Series in Data Management Systems. Morgan Kaufmann Publishers Inc., 1997.

[bhattacharyya1977a] Gouri K. Bhattacharyya and Richard A. Johnson. Statistical Concepts and Methods. John Wiley & Sons, 1977.

[biggs1989a] Norman L. Biggs. Discrete Mathematics. Oxford University Press, revised edition, 1989.

[bird1988a] Richard Bird and Philip Wadler. Introduction to Functional Programming. Prentice Hall International Series in Computer Science. Prentice Hall, 1988.

[bishop2001a] Judith Bishop. Java Gently. International Computer Science Series. Addison-Wesley, 3 edition, 2001.

[bjornestad1994a] Harald Bjørnestad, Ulf Henning Olsson, Svein Søyland, and Frank Tolcsiner. Matematikk for Økonomi og Samfunnsfag. HøyskoleForlaget, 2 edition, 1994.

[bjornestad1994b] Harald Bjørnestad, Ulf Henning Olsson, Svein Søyland, and Frank Tolcsiner. Matematikk for Økonomi og Samfunnsfag — Fasit til 2. utgave. HøyskoleForlaget, 1994.

[blair1998a] Gordon S. Blair and Jean-Bernard Stefani. Open Distributed Processing and Multi-

media. Addison-Wesley, 1998.

[blair1998b] Gordon S. Blair, Lynne Blair, Howard Bowman, and Amanda Chetwynd. Formal

Specification of Distributed Multimedia Systems. UCL Press, 1998.

[blair1998f] John D. Blair. SAMBA, Integrating UNIX and Windows. Specialized System Con-

sultants, Inc., 1998.

[blakley1999a] Bob Blakley. CORBA Security. Object Technology Series. Addison-Wesley, 1999. [bolinger1995a] Don Bolinger and Tan Bronson. Applying RCS and SCCS. A Nutshell Handbook.

O’Reilly & Associates, Inc., 1995.

[boney2002a] James Boney. Cisco IOS in a Nutshell. A Nutshell Handbook. O’Reilly & Associates,

Inc., 2002.

[bowman2001a] Howard Bowman and John Derrick, editors. Formal Methods for Distributed

Processing. Cambridge University Press, 2001.

[box1998a] Don Box. Essential COM. Object Technology Series. Addison-Wesley, 1998.

[box2000a] Don Box, Aaron Skonnard, and John Lam. Essential XML, Beyond Markup. The DevelopMentor Series. Addison-Wesley, 2000.

[bradley2000a] Neil Bradley. The XSL Companion. Addison-Wesley, 2000. [brazier1994a] F. M. T. Brazier and D. Johansen, editors. Distributed Open Systems. IEEE

Computer Society Press, 1994.

[brinch-hansen1973a] Per Brinch Hansen. Operating System Principles. Prentice Hall Series in

Automatic Computation. Prentice Hall, 1973.

[bringhurst1996] Robert Bringhurst. The Elements of Typographic Style. Hartley & Marks, Pub-

lishers, 2 edition, 1996.

[britton2001a] Chris Britton. IT Architectures and Middleware. Addison-Wesley, 2001. [brown1995a] C. Wayne Brown and Barry J. Shepherd. Graphics File Formats. Prentice Hall,

1995.

[brown1998a] William J. Brown, Raphael C. Malveau, Hays W. “Skip” McCormick III, and Thomas J. Mowbray. AntiPatterns: Refactoring Software, Architectures, and Projects in Crisis. John Wiley & Sons, 1998.

[brown2000a] Alan W. Brown. Large-Scale Component-Based Development. Object and Compo-

nent Technology Series. Prentice Hall, 2000.

[buford1994a] John F. Koegel Buford, editor. Multimedia Systems. Addison-Wesley/ACM Press,

1994.

[campbell1987a] Joe Campbell. C Programmer’s Guide to Serial Communications. SAMS, 1992. [campbell1997a] Andrew Campbell and Klara Nahrstedt, editors. Building QoS into Distributed

Systems. IFIP/Chapman & Hall, 1997. IWQOS’97.

[cerami2002a] Ethan Cerami. Web Services Essentials. A Nutshell Handbook. O’Reilly & Asso-

ciates, Inc., 2002.

[chang2000a] Ben Chang, Mark Scardina, K. Karun, Stefan Kiritzov, Ian Macky, Anguel Novoselsky, and Niranjan Ramakrishnan. Oracle XML Handbook. Oracle Press/Osborne McGraw-Hill, 2000.

[chapman2002a] Nigel Chapman and Jenny Chapman. Digital Media Tools. John Wiley & Sons, 2002.

[cheesman2001a] John Cheesman. UML Components. The Component Software Series. Addison-Wesley, 2001.

[cheswick1994a] William R. Cheswick and Steven M. Bellovin. Firewalls and Internet Security. Addison-Wesley Professional Computing Series. Addison-Wesley, 1994.

[cohen1991a] Daniel I. A. Cohen. Introduction to Computer Theory. John Wiley & Sons, revised edition, 1991.

[comer1999a] Douglas E. Comer. Computer Networks and Internets. Prentice Hall, 2 edition, 1999.

[comer2000a] Douglas E. Comer. Internetworking with TCP/IP, Principles, Protocols, and Architectures, volume 1. Prentice Hall, 4 edition, 2000.

[comer2001a] Douglas E. Comer. Computer Networks and Internets. Prentice Hall, 3 edition, 2001.

[conard2000a] James Conard, Patrick Dengler, Brian Francis, Jay Glynn, Burton Harvey, Billy Hollis, Rama Ramachandran, John Schenken, Scott Short, and Chris Ullman. Introducing .NET. Wrox Press Ltd, 2000.

[cormen1990a] Thomas H. Cormen, Charles E. Leiserson, and Ronald L. Rivest. Introduction to Algorithms. The MIT Electrical Engineering and Computer Science Series. The MIT Press/McGraw-Hill Book Company, 1990.

[cormen2001a] Thomas H. Cormen, Charles E. Leiserson, Ronald L. Rivest, and Clifford Stein. Introduction to Algorithms. The MIT Press/McGraw-Hill Book Company, 2 edition, 2001.

[costales1997a] Bryan Costales and Eric Allman. Sendmail. A Nutshell Handbook. O’Reilly & Associates, Inc., 2 edition, 1997.

[coulouris1988a] George F. Coulouris and Jean Dollimore. Distributed Systems — Concepts and Design. International Computer Science Series. Addison-Wesley, 1988.

[coulouris1994a] George Coulouris, Jean Dollimore, and Tim Kindberg. Distributed Systems — Concepts and Design. International Computer Science Series. Addison-Wesley, 2 edition, 1994.

[coulouris2001a] George Coulouris, Jean Dollimore, and Tim Kindberg. Distributed Systems — Concepts and Design. International Computer Science Series. aw, 3 edition, 2001.

[crichlow1988a] Joel M. Crichlow. An Introduction to Distributed and Parallel Computing. Prentice Hall, 1988.

[crowcroft1995a] Jon Crowcroft. Open Distributed Systems. Artech House, 1995.

[curry1992a] David A. Curry. UNIX System Security — A Guide for Users and System Administrators. Addison-Wesley Professional Computing Series. Addison-Wesley, 1992.

[cypser1978a] R. J. Cypser. Communications Architecture for Distributed Systems. The Systems Programming Series. Addison-Wesley, 1978.

[dahl1972a] O.-J. Dahl, E. W. Dijkstra, and C. A. R. Hoare. Structured Programming. Academic Press, 1972.

[dahl1992a] Ole-Johan Dahl. Verifiable Programming. Prentice Hall International Series in Com-

puter Science. Prentice Hall, 1992.

[dalheimer1999a] Matthias Kalle Dalheimer. Programming with Qt. A Nutshell Handbook.

O’Reilly & Associates, Inc., 1999.

[date1995a] C. J. Date. An Introduction to Database Systems. Addison-Wesley Systems Program-

ming Series. Addison-Wesley, 4 edition, 1995.

[davie2000a] Bruce Davie and Yakov Rekhter. MPLS Technology and Applications. Morgan Kauf-

mann Publishers Inc., 2000.

[de-blasi1990a] Mario De Blasi. Computer Architecture. International Computer Science Series.

Addison-Wesley, 1990.

[dibona1999a] Chris DiBona, Sam Ockman, and Mark Stone, editors. Open Sources: Voices of

the Open Source Revolution. O’Reilly & Associates, Inc., 1999.

[dick2000a] Kevin Dick. XML, A Managers’s Guide. Addison-Wesley Information Technology

Series. Addison-Wesley, 2000.

[dijkstra1976a] Edsger W. Dijkstra. A Discipline of Programming. Prentice Hall Series in Auto-

matic Computation. Prentice Hall, 1976.

[dijkstra1982a] Edsger W. Dijkstra. Selected Writings on Computing: A personal Perspective.

Texts and Monographs in Computer Science. Springer-Verlag, 1982.

[dijkstra1988a] Edsger W. Dijkstra and W. H. J. Feijen. A Method of Programming. Addison-

Wesley, 1988.

[dijkstra1990a] Edsger W. Dijkstra, editor. Formal Development of Programs and Proofs. Uni-

versity of Texas at Austin Year of Programming Series. Addison-Wesley, 1990.

[dougherty1990a] Dale Dougherty. Sed & Awk. A Nutshell Handbook. O’Reilly & Associates,

Inc., 1990.

[dowding1988a] Barry Dowding. Principles of Electronics. Prentice Hall, 1988. [dsouza1999a] Desmond F. D’Souza and Alan Cameron Wills. Objects, Components, and Frame-

works with UML. Object Technology Series. Addison-Wesley, 1999.

[dubois1993a] Paul DuBois. Software Portability with imake. A Nutshell Handbook. O’Reilly &

Associates, Inc., 1993.

[edwards1999a] W. Keith Edwards. Core Jini. The Sun Microsystem Press Java Series. Prentice

Hall PTR, 1999.

[elmasri1994a] Ramez Elmasri and Shamkant B. Navathe. Fundamentals of Database Systems.

Benjamin/Cummings, 2 edition, 1994.

[emmerich2000a] Wolfgang Emmerich. Engineering Distributed Objects. John Wiley & Sons, 2000. [esakov1989a] Jeffrey Esakov and Tom Weiss. Data Structures — An Advanced Approach Using

C. Prentice Hall, 1989.

[flanagan1996a] David Flanagan. Java in a nutshell. A Nutshell Handbook. O’Reilly & Associates,

Inc., 1996.

[fluckiger1995a] François Fluckiger. Understanding Networked Multimedia — Applications and Technology. Prentice Hall, 1995.

[forouzan2000a] Behrouz A. Forouzan. TCP/IP Protocol Suite. Computer Science Series.

McGraw-Hill Book Company, 2000.

[forouzan2001a] Behrouz A. Forouzan. Data Communication and Networking. Computer Science

Series. McGraw-Hill Book Company, 2 edition, 2001.

[fowler2000a] Martin Fowler and Kendall Scott. UML Distilled. ot. aw, 2 edition, 2000. [fraleigh1989a] John B. Fraleigh. A First Course in Abstract Algebra. Addison-Wesley, 4 edition,

1989.

[freund1992a] John E. Freund. Mathematical Statistics. Prentice Hall, 5 edition, 1992. [frolund1996a] Svend Frølund. Coordinating Distributed Objects: An Actor Based Approach to

Synchronization. The MIT Press, 1996.

[fung2001a] Khun Yee Fung. XSLT, Working with XML and HTML. Addison-Wesley, 2001. [furth1995a] Borko Furth and Milan Milenkovic, editors. A Guided Tour of Multimedia Systems

and Applications. IEEE Computer Society Press, 1995.

[furth1995b] Borko Furth, Stephen W. Smoliar, and Hongjiang Zhang. Video and Image Processing in Multimedia Systems. The Kluwer International Series in Engineering and Computer Science. Kluwer Academic Publishers, 1995.

[gallier1987a] Jean H. Gallier. Logic for Computer Science. Harper & Row Computer Science and

Technology Series. John Wiley & Sons, 1987.

[gamma95a] Erich Gamma, Richard Helm, Ralph Johnson, and John Vlissides. Design Patterns: Elements of Reusable Object-Oriented Software. Addison-Wesley Professional Computing Series. Addison-Wesley, 1995.

[garg1996a] Vijay K. Garg. Principles of Distributed Systems. Kluwer Academic Publishers, 1996. [gehani1988a] Narain Gehani and Andrew D. McGettrick, editors. Concurrent Programming.

International Computer Science Series. Addison-Wesley, 1988.

[geraghty1999a] Ronan Geraghty, Sean Joyce, Tom Moriarty, and Gary Noone. COM-CORBA

Interoperability. Prentice Hall Series on Microsoft Technologies. Prentice Hall, 1999.

[gibbs1995a] Simon J. Gibbs and Dionysios C. Tsichritzis. Multimedia Programming — Objects,

Environments and Frameworks. Addison-Wesley/ACM Press, 1995.

[girdley2002a] Michael Girdley, Rob Woollen, and Sandra L. Emerson. J2EE Applications and

BEA WebLogic Server. Prentice Hall, 2002.

[glickenstein1997a] Bob Glickenstein. GNU Emacs Extension. A Nutshell Handbook. O’Reilly &

Associates, Inc., 1997.

[goldfarb2000a] Charles F. Goldfarb and Paul Prescod. The XML handbook. The Charles F.

Goldfarb Series on Open Information Management. Prentice Hall, 2 edition, 2000.

[gomez2000a] Paco Gómez and Peter Zadrozny. Java 2 Enterprise Edition with BEA WebLogic

Server. Wrox Press Ltd, 2000.

[gonnet1984a] G. H. Gonnet. Handbook of Algorithms and Data Structures. International Com-

puter Science Series. Addison-Wesley, 1984.

[goodhearth1994a] Berny Goodheart and James Cox. The Magic Garden Explained — The Internals of UNIX System V Release 4, An Open System Design. Prentice Hall, 1994.

[goodrich2001] Michael T. Goodrich and Roberto Tamassia. Data Structures and Algorithms in

Java. John Wiley & Sons, 2 edition, 2001.

[goossens1994a] Michel Goossens, Frank Mittelbach, and Alexander Samarin. The L^{A}TEX Com-

panion. Addison-Wesley, 1994.

[goossens1997a] Michel Goossens, Sebastian Rahtz, and Frank Mittelbach. The L^{A}TEX Graphics Companion. Addison-Wesley Series on Tools and Techniques for Computer Typesetting. Addison-Wesley, 1997.

[goossens1999a] Michel Goossens and Sebastian Rahtz. The L^{A}TEX Web Companion. Addison-Wesley Series on Tools and Techniques for Computer Typesetting. Addison-Wesley, 1999.

[goscinski1991a] A. Goscinski. Distributed Operating Systems — The Logical Design. Addison-

Wesley, 1991.

[graham1996a] Paul Graham. ANSI Common Lisp. Prentice Hall Series in Artificial Intelligence.

Prentice Hall, 1996.

[graham2002a] Steve Graham, Simeon Simeonov, Toufic Boubez, Doug Davis, Glen Daniels,

Yuichi Nakamura, and Ryo Neyama. Building Web Services with Java. SAMS, 2002.

[grayson2000a] John E. Grayson. Python and Tkinter Programming. Manning, 2000. [grimaldi1985a] Ralph P. Grimaldi. Discrete and Combinatorial Mathematics — An Applied In-

troduction. Addison-Wesley, 1985.

[grosso2002a] William Grosso. Java RMI. O’Reilly & Associates, Inc., 2001. [habermann1983a] A. Nico Habermann and Dewayne E. Perry. Ada for Experienced Programmers.

Addison-Wesley Series in Computer Science. Addison-Wesley, 1983.

[hacker1999a] Scot Hacker. The BeOS Bible. Peachpit Press, 1999. [hahn1993a] Harley Hahn. A Student’s Guide to UNIX. Computer Science Series. McGraw-Hill

Book Company, 1993.

[hahn1994a] Harley Hahn. Open Computing — Unix Unbound. Osborne McGraw-Hill, 1994. [halsall2001a] Fred Halsall. Multimedia Communications. Addison-Wesley, 2001. [halter2001a] Steven L. Halter and Steven J. Munroe. Enterprise Java Performance. Java 2

Platform, Enterprise Edition Series. Prentice Hall/Sun Microsystem, 2001.

[hammel2001a] Michael J. Hammel. Essential GIMP for Web Professionals. Prentice Hall, 2001. [hammond2000a] Mark Hammond and Andy Robinson. Python Programming on Win32. A Nut-

shell Handbook. O’Reilly & Associates, Inc., 2000.

[harbison1987a] Samuel P. Harbison and Guy L. Steele, Jr. C: A Reference Manual. Prentice Hall

Software Series. Prentice Hall, 2 edition, 1987.

[harel1992] David Harel. Algorithmics. Addison-Wesley, 2 edition, 1992. [haviland1987a] Keith Haviland and Ben Salama. UNIX System Programming. International

Computer Science Series. Addison-Wesley, 1987.

[helfrick1990a] Albert D. Helfrick and William D. Cooper. Modern Electronic Instrumentation and Measurement Techniques. Prentice Hall, 1990.

[hellmark2000a] Christer Hellmark. Typorafisk Håndbok. Spartacus, 2000. [hennessy1990a] John L. Hennessy and David A. Patterson. Computer Architecture — A Quan-

titative Approach. Morgan Kaufmann Publishers Inc., 1990.

[henning1999a] Michi Henning and Steve Vinoski. Advanced CORBA Programming with C++.

Addison-Wesley Professional Computing Series. Addison-Wesley, 1999.

[henshall1988a] John Henshall and Sandy Shaw. OSI Explained — End-to-end Computer Communication Standards. Computer Communications and Networking. Ellis Horwood Limited, 1988.

[hinchey1995a] Michael G. Hinchey and Jonathan P. Bowen, editors. Applications of Formal

Methods. Prentice Hall International Series in Computer Science. Prentice Hall, 1995.

[hoare1985a] C. A. R. Hoare. Communicating Sequential Processes. Prentice Hall International

Series in Computer Science. Prentice Hall, 1985.

[hoare1989a] C. A. R. Hoare and C. B. Jones. Essays in Computing Science. Prentice Hall

International Series in Computer Science. Prentice Hall, 1989.

[hodges1977a] Wilfrid Hodges. Logic — An Introduction to Elementary Logic. Penguin Books,

1977.

[hoenig1998a] Alan Hoenig. TEX Unbound: L^{A}TEX & TEX Stratergies for Fonst, Graphics, & More.

Oxford University Press, 1998.

[horstmann2000a] Cay Horstmann. Computer Concepts with Java 2 Essentials. John Wiley &

Sons, 2 edition, 2000.

[huitema1998a] Christian Huitema. IPv6, The New Internet Protocol. Prentice Hall, 1998. [huitema2000a] Christian Huitema. Routing in the Internet. Prentice Hall PTR, 2 edition, 2000. [hunt1998a] Craig Hunt. TCP/IP Network Administration. A Nutshell Handbook. O’Reilly &

Associates, Inc., 2 edition, 1998.

[hunt2000a] Andrew Hunt and David Thomas. The Pragmatic Programmer. Addison-Wesley,

2000.

[istad1992a] Roy M. Istad. Forkurs i Matematikk. Bedriftøkonomens Forlag, 1992. [jaeschke1986a] Rex Jaeschke. Solutions in C. Addison-Wesley, 1986. [jones1989a] Oliver Jones. Introduction to the X Window System. Prentice Hall, 1989. [jones2002a] Christopher A. Jones and Jr. Fred L. Drake. Python & XML. A Nutshell Handbook.

O’Reilly & Associates, Inc., 2002.

[jubin2000a] Henri Jubin, Jürgen Friedrichs, and the Jalapeño Team. Enterprise JavaBeans by

Example. Prentice Hall PTR/IBM, 2000.

[karp2002a] David A. Karp, Tim O’Reilly, and Troy Mott. Windows XP in a nutshell. A Nutshell

Handbook. O’Reilly & Associates, Inc., 2002.

[kassem2000a] Nicholas Kassem et al. Designing Enterprise Applications with the Java 2 Platform, Enterprise Edition. The Java Series, Enterprise Edition. Addison-Wesley/Sun Microsystem, 2000.

[kavi1992a] Krishna M. Kavi, editor. Real-Time Systems — Abstractions, Languages, and Design Methodologies. IEEE Computer Society Press Tutorial. IEEE Computer Society Press, 1992.

[kay2000a] Michael Kay. XSLT Programmers’s Reference. Wrox Press Ltd, 2000.

[kay2001a] Michael Kay. XSLT Programmers’s Reference. Wrox Press Ltd, 2 edition, 2001. [keogh1994a] James Keogh and Remon Lapid. Open Computing — Guide to the Best Free UNIX

Utilities. Osborne McGraw-Hill, 1994.

[kernighan1988a] Brian W. Kernighan and Dennis M. Ritchie. The C Programming Language.

Prentice Hall, 2 edition, 1988.

[kernighan1999a] Brian W. Kernighan and Rob Pike. The Practice of Programming. Addison-

Wesley Professional Computing Series. Addison-Wesley, 1999.

[kiczales1991a] Gregor Kiczales, Jim des Rivières, and Daniel G. Bobrow. The Art of the Metaob-

ject Protocol. The MIT Press, 1991.

[knuth1984a] Donald E. Knuth. The TEXBook, volume A of Computers & Typesetting. Addison-

Wesley, 1984.

[knuth1986a] Donald E. Knuth. The Book, volume C of Computers & Typesetting.

Addison-Wesley, 1986.

[knuth1993a] Donald E. Knuth. The Stanford GraphBase — A Platform for Combinatorial Com-

puting. ACM Press/Addison-Wesley, 1993.

[knuth1997a] Donald E. Knuth. The Art of Computer Programming — Fundamental Algorithms,

volume 1. Addison-Wesley, 3 edition, 1997.

[knuth1998a] Donald E. Knuth. The Art of Computer Programming — Seminumerical Algorithms,

volume 2. Addison-Wesley, 3 edition, 1998.

[knuth1998b] Donald E. Knuth. The Art of Computer Programming — Sorting and Searching,

volume 3. Addison-Wesley, 3 edition, 1997.

[knuth1999a] Donald E. Knuth. Digital Typography. Number 78 in CSLI Lecture Notes. CSLI

Publications, Stanford, California, 1999.

[knuth2000a] Donald E. Knuth. Selected Papers on Analysis of Algorithms. CSLI, 2000. [kopka1995a] Helmut Kopka and Patrick W. Daly. A Guide to L^{A}TEX2e — Document Preparation

for Beginners and Advanced Users. Addison-Wesley, 2 edition, 1995.

[krol1992a] Ed Krol. The Whole Internet: User’s Guide & Catalog. A Nutshell Handbook.

O’Reilly & Associates, Inc., 1992.

[kurose2000a] James F. Kurose and Keith W. Ross. Computer Networking. Addison-Wesley, 2000. [lamport1986e] Leslie Lamport. L^{A}TEX, A Document Preparation System — User’s Guide & Ref-

erence Manual. Addison-Wesley, 1986.

[lamport1994a] Leslie Lamport. L^{A}TEX, A Document Preparation System — User’s Guide & Ref-

erence Manual. Addison-Wesley, 2 edition, 1994.

[laplante1993a] Phillip A. Laplante. Real-Time Systems Design and Analysis — An Engineer’s

Handbook. IEEE Press/IEEE Computer Society Press, 1993.

[latteier2002a] Amos Latteier and Michael Pelletier. The Zope Book. New Riders, 2002. [leavens2000a] Gary T. Leavens and Murali Sitaraman, editors. Foundations of Component-Based

Systems. Cambridge University Press, 2000.

[leffler1989a] Samuel J. Leffler, Marshall Kirk McKusick, Michael J. Karels, and John S. Quarterman. The Design and Implementation of the 4.3BSD UNIX Operating System. Addison-Wesley Series in Computer Science. Addison-Wesley, 1989.

[leffler1991a] Samuel J. Leffler and Marshall Kirk McKusick. The Design and Implementation of the 4.3BSD UNIX Operating System: Answer Book. Addison-Wesley Series in Computer Science. Addison-Wesley, 1991.

[lerdorf2000a] Rasmus Lerdorf. PHP Pocket Reference. A Nutshell Handbook. O’Reilly & Asso-

ciates, Inc., 2000.

[lervik2000a] Else Lervik and Vegard B. Havdal. Programmering i Java. Gyldendal Akademisk,

2000.

[levy1994a] Steven Levy. Hackers. Penguin Books, 1994. First published by Dell Publishing,

1984.

[libes1989a] Don Libes and Sandy Ressler. Life with UNIX — A Guide for Everyone. Prentice

Hall, 1989.

[linden1994a] Peter van der Linden. Expert C Programming — Deep C Secrets. SunSoft Press.

Prentice Hall, 1994.

[lippman1991a] Stanley B. Lippman. C++ Primer. Addison-Wesley, 2 edition, 1991.

[lu1996a] Guojun Lu. Communication and Computing for Distributed Multimedia Systems. Artech House, 1996.

[lundh2000a] Fredirk Lundh. The Standard Python Library. PythonWare, 2000. [lutz1996a] Mark Lutz. Programming Python. A Nutshell Handbook. O’Reilly & Associates, Inc.,

1996.

[lutz1998a] Mark Lutz. Python Pocket Reference. A Nutshell Handbook. O’Reilly & Associates,

Inc., 1998.

[lutz1999a] Mark Lutz and David Ascher. Learning Python. A Nutshell Handbook. O’Reilly &

Associates, Inc., 1999.

[lutz2001a] Mark Lutz. Programming Python. A Nutshell Handbook. O’Reilly & Associates, Inc.,

2 edition, 2001.

[lynch1994a] Nancy Lynch, Michael Merritt, William Weihl, and Alan Fekete. Atomic Transac-

tions. Morgan Kaufmann Publishers Inc., 1994.

[lynch1996a] Nancy A. Lynch. Distributed Algorithms. Morgan Kaufmann Publishers Inc., 1996. [maartmann-moe1991a] Erling Maartmann-Moe. Multimedia. Universitetsforlaget, 1991. [maekawa1987a] Mamoru Maekawa, Arthur E. Oldehoeft, and Rodney R. Oldehoeft. Operating

Systems — Advanced Concepts. Benjamin/Cummings, 1987.

[malmstadt1981a] Howard V. Malmstadt, Christie G. Enke, and Stanley R. Crouch. Electronics

and Instrumentation for Scientists. Benjamin/Cummings, 1981.

[marinescu2002a] Floyd Marinescu. EJB Design Patterns. John Wiley & Sons, 2002. [markel1994a] Mike Markel. Writing in the Technical Fields — A Step-by-Step Guide for Engi-

neers, Scientists, and Technicians. IEEE Press, 1994.

[marsden1985a] Jerrold Marsden and Alan Weinstein. Calculus, volume I. Springer-Verlag, 1985.

[marsden1985b] Jerrold Marsden and Alan Weinstein. Calculus, volume II. Springer-Verlag, 1985. [marsden1985c] Jerrold Marsden and Alan Weinstein. Calculus, volume III. Springer-Verlag, 1985. [martin1986a] Johannes J. Martin. Data Types and Data Structures. Prentice Hall International

Series in Computer Science. Prentice Hall, 1986.

[martin1988a] James Martin. Data Communication Technology. Prentice Hall, 1988. [matena2001a] Vlada Matena and Beth Stearns. Applying Enterprise JavaBeans. The Java Series,

Enterprise Edition. Addison-Wesley/Sun Microsystem, 2001.

[mazza1994a] C. Mazza, J. Fairclough, B. Melton, D. de Pablo, A. Scheffer, and R. Stevens.

Software Engineering Standards. Prentice Hall, 1994.

[mcconnell2001a] Jeffrey J. McConnell. Analysis of Algorithms: An Active Learning Approach.

Jones and Bartlett Computer Science. Jones and Bartlett Publishers, 2001.

[mcgrath2000a] Sean Mcgrath. XML Processing with Python. The Charles F. Goldfarb Series on

Open Information Management. Prentice Hall PTR, 2000.

[mcmillan1998a] Michael McMillan. Perl from the Ground Up. Osborne McGraw-Hill, 1998. [meyer2000a] Eric A. Meyer. Cascading Style Sheets, The Definitive Guide. A Nutshell Handbook.

O’Reilly & Associates, Inc., 2000.

[meyers1992a] Scott Meyers. Effective C++ — 50 Specific Ways to Improve Your Programs and

Designs. Addison-Wesley Professional Computing Series. Addison-Wesley, 1992.

[milenkovic1992a] Milan Milenković. Operating Systems — Concepts and Design. Computer Sci-

ence Series. McGraw-Hill Book Company, 2 edition, 1992.

[milner1989a] Robin Milner. Communication and Concurrency. Prentice Hall International Series

in Computer Science. Prentice Hall, 1989.

[milner1989b] Robin Milner. Communication and Concurrency — Solutions Manual. Prentice

Hall International Series in Computer Science. Prentice Hall, 1989.

[minsky1988a] Marvin Minsky. The Society of Mind. Touchstone, Simon & Schuster, 1988. [monson-haefel2000a] Richard Monson-Haefel. Enterprise JavaBeans. The Java Series. O’Reilly

& Associates, Inc., 2 edition, 2000.

[morrison2000a] Michael Morrison. XML Unleashed. SAMS, 2000. [mowbray1997a] Thomas J. Mowbray and Raphael C. Malveau. CORBA Design Patterns. John

Wiley & Sons, 1997.

[mowbray1997b] Thomas J. Mowbray and William Ruh. Inside CORBA: Distributed Object Stan-

dards and Applications. Object Technology Series. Addison-Wesley, 1997.

[mui1994a] Linda Mui and Valerie Quercia. X User Tools. O’Reilly & Associates, Inc., 1994. [mullender1989a] Sape Mullender, editor. Distributed Systems. ACM Press Frontier Series. ACM

Press/Addison-Wesley, 1989.

[mullender1993a] Sape Mullender, editor. Distributed Systems. ACM Press Frontier Series.

Addison-Wesley, 2 edition, 1993.

[mullin1989a] Mark Mullin. Object Oriented Program Design — With Examples in C++. Addison-Wesley, 1989.

[myerson2002a] Judith M. Myerson. The Complete Book of Middleware. Auerbach, 2002. [naur1992a] Peter Naur. Computing: A Human Activity. ACM Press. Addison-Wesley, 1992. [nemeth1995a] Evi Nemeth, Garth Snyder, Scott Seebass, and Trent R. Hein. Unix System Ad-

ministration Handbook. Prentice Hall, 2 edition, 1995.

[neumann1995a] Peter G. Neumann. Computer Related Risks. ACM Press/Addison-Wesley, 1995. [newham1995a] Cameron Newham and Bill Rosenblatt. Learning the Bash shell. A Nutshell

Handbook. O’Reilly & Associates, Inc., 1995.

[nichols1996a] Bradford Nichols, Dick Buttlar, and Jacqueline Proulx Farrell. Pthreads Program-

ming. A Nutshell Handbook. O’Reilly & Associates, Inc., 1996.

[nicholson1995a] W. Keith Nicholson. Linear Algebra with Applications. PWS Publishing Com-

pany, 3 edition, 1995.

[norvig1992a] Peter Norvig. Paradigms of Artificial Intelligence Programming: Case studies in

Common Lisp. Morgan Kaufmann Publishers Inc., 1992.

[oestereich1999a] Bernd Oestereich. Developing Software with UML. Object Technology Series.

Addison-Wesley, 1999.

[oram1991a] Andrew Oram and Steve Talbott. Managing Projects with Make. A Nutshell Hand-

book. O’Reilly & Associates, Inc., 2 edition, 1991.

[orfali1997a] Robert Orfali and Dan Harkey. Client/Server Programming with Java and CORBA.

John Wiley & Sons, 1997.

[orfali1998a] Robert Orfali and Dan Harkey. Client/Server Programming with Java and CORBA.

John Wiley & Sons, 2 edition, 1998.

[otte1996a] Randy Otte, Paul Patrick, and Mark Roy. Understanding CORBA. Prentice Hall,

1996.

[ousterhout1994a] John K. Ousterhout. Tcl and the Tk Toolkit. Addison-Wesley Professional

Computing Series. Addison-Wesley, 1994.

[papadimitriou1994a] Christos H. Papadimitriou. Computational Complexity. Addison-Wesley,

1994.

[peek1993a] Jerry Peek, Tim O’Reilly, and Mike Loukides. UNIX Power Tools. A Nutshell

Handbook. O’Reilly & Associates, Inc./Bantam Book, 1993.

[peek1995a] Jerry Peek. MH & xmh — Email for Users & Programmers. A Nutshell Handbook.

O’Reilly & Associates, Inc., 3 edition, 1995.

[perlman2000a] Radia Perlman. Interconnections. Addison-Wesley Professional Computing Series.

Addison-Wesley, 2 edition, 2000.

[perrone2000a] Paul J. Perrone and Venkata S. R. “Krishna” R. Chaganti. Building Java Enter-

prise Systems with j2EE. SAMS, 2000.

[peterson2000a] Larry L. Peterson and Bruce S. Davie. Computer Networks, A System Approach. The Morgan Kaufmann Series in Networking. Morgan Kaufmann Publishers Inc., 2 edition, 2000.

[platt1999a] David S. Platt. Understanding COM+. Developer Technology Series. Microsoft Press, 1999.

[pohl2000a] Ira Pohl and Charlie McDowell. Java by Dissection. Addison-Wesley, 2000. [pope1998a] Alan Pope. The CORBA Reference Guide. Addison-Wesley, 1998. [pritchard1999a] Jason Pritchard. COM and CORBA Side by Side. Addison-Wesley, 1999. [purdy2000a] Gregor N. Purdy. CVS Pocket Reference. A Nutshell Handbook. O’Reilly & Asso-

ciates, Inc., 2000.

[raghavan1998a] S. V. Raghavan and Satish K. Tripathi. Networked Multimedia Systems: Con-

cepts, Architecture, and Design. Prentice Hall, 1998.

[ray2001a] Erik T. Ray. Learning XML. A Nutshell Handbook. O’Reilly & Associates, Inc., 2001.

[reade1989a] Chris Reade. Elements of Functional Programming. International Computer Science

Series. Addison-Wesley, 1989.

[reese2002a] George Reese, Randy Jay Yarger, and Tim King. Managing and Using MySQL. A

Nutshell Handbook. O’Reilly & Associates, Inc., 2 edition, 2002.

[rhodes2002a] Neil Rhodes and Julie McKeehan. Palm OS Programming: The Developer’s Guide.

A Nutshell Handbook. O’Reilly & Associates, Inc., 2 edition, 2002.

[risnes1992a] Martin Risnes. Matematikk med anvendelse i økonomi. Universitetsforlaget, 4 edi-

tion, 1992.

[roman1999c] Ed Roman. Mastering Enterprise JavaBeans and the Java 2 Platform, Enterprise

Edition. Wiley Computer Publishing. John Wiley & Sons, 1999.

[roman2002a] Ed Roman, Scott Ambler, and Tyler Jewell. Mastering Enterprise JavaBeans. John

Wiley & Sons, 2 edition, 2002.

[roscoe1994a] A. W. Roscoe, editor. A Classical Mind — Essays in the Honour of C. A. R. Hoare.

Prentice Hall International Series in Computer Science. Prentice Hall, 1994.

[rose1990a] Marshall T. Rose. The Open Book — A Practical Perspective on OSI. Prentice Hall,

1990.

[rosen1988a] Kenneth H. Rosen. Elementary Number Theory and its Application. Addison-Wesley,

2 edition, 1988.

[rosenberger1998a] Jeremy Rosenberger. Sams’ Teach Yourself CORBA in 14 Days. SAMS, 1998. [rubin1999a] William Rubin and Marshall Brain. Understanding DCOM. Prentice Hall Series on

Microsoft Technologies. Prentice Hall, 1999.

[ruh2000a] William Ruh, Thomas Herron, and Paul Klinker. IIOP Complete: Understanding CORBA and Middleware Interoperability. Object Technology Series. Addison-Wesley, 2000.

[salus1994a] Peter H. Salus. A Quarter Centery of UNIX. Addison-Wesley, 1994. [santifaller1991a] Michael Santifaller. TCP/IP and NFS — Internetworking in a UNIX Environ-

ment. Addison-Wesley, 1991.

[savant1987a] C. J. Savant, Jr., Martin S. Roden, and Gordon L. Carpenter. Electronic Circuit

Design — An Engineering Approach. Benjamin/Cummings, 1987.

[schildt1987a] Herbert Schildt. C: The Complete Reference. Osborne McGraw-Hill, 1987.

[schmidt2000a] Douglas Schmidt, Michael Stal, Hans Rohnert, and Frank Buschmann. Pattern-Oriented Software Architecture, Patterns for Concurrent and Networked Objects, volume 2 of Wiley Series in Software Design Patterns. John Wiley & Sons, 2000.

[schneier1996b] Bruce Schneier. Applied Cryptography. John Wiley & Sons, 2 edition, 1996. [schoonover1992a] Michael A. Schoonover, John S. Bowie, and William R. Arnold. GNU Emacs —

UNIX Text Editing and Programming. Hewlett-Packard Press Series. Addison-Wesley, 1992.

[schwaderer1985a] W. David Schwaderer. C Wizard’s Programming Reference. Wiley Press Books.

John Wiley & Sons, 1985.

[scribner2000a] Kennard Scribner and Mark C. Stiver. Understanding SOAP. SAMS, 2000. [serain1999a] Daniel Serain. Middleware. Practitioner Series. Springer-Verlag, 1999. [sessions2000a] Roger Sessions. COM+ and the Battle for the Middle Tier. Wiley Computer

Publishing. John Wiley & Sons, 2000.

[shakespeare1999a] Norman Shakespeare. Countdown to the Year 2000. A Nutshell Handbook.

O’Reilly & Associates, Inc., 1999.

[shatza1989a] Sol M. Shatz and Jia-Ping Wang, editors. Tutorial: Distributed-Software Engineer-

ing. IEEE Computer Society Press, 1989.

[siegel1996a] Jon Siegel. CORBA — Fundamentals and Programming. John Wiley & Sons, 1996. [siegel2000a] Jon Siegel. CORBA 3 — Fundamentals and Programming. Wiley Computer Pub-

lishing. John Wiley & Sons/OMG Press, 2 edition, 2000.

[silberschatz1991a] A. Silberschatz, J. Peterson, and P. Galvin. Operating System Concepts.

Addison-Wesley, 3 edition, 1991.

[silberschatz1994a] A. Silberschatz and P. Galvin. Operating System Concepts. Addison-Wesley,

4 edition, 1994.

[silberschatz1994b] A. Silberschatz and P. Galvin. Instructor’s Manual: Operating System Con-

cepts. Addison-Wesley, 4 edition, 1994.

[simpson2001a] John E. Simpson. Just XML. Prentice Hall, 2 edition, 2001. [skonnard2002a] Aaron Skonnard and Martin Gudgin. Essential XML Quick Reference. The

DevelopMentor Series. Addison-Wesley, 2002.

[slama1999a] Dirk Slama, Jason Garbis, and Perry Russell. Enterprise CORBA. Prentice Hall,

1999.

[snell2002a] James Snell, Doug Tidwell, and Pavel Kulchenko. Programming Web Services with

SOAP. A Nutshell Handbook. O’Reilly & Associates, Inc., 2002.

[sobell1995a] Mark G. Sobell. UNIX System V — A Practical Guide. Benjamin/Cummings, 3

edition, 1995.

[sobell1995b] Mark G. Sobell. Hands-On UNIX — A Practical Guide with the Essentials of CO-

HERENT. Benjamin/Cummings, 1995.

[son1995a] Sang H. Son, editor. Advances in Real-Time Systems. Prentice Hall, 1995. [sperschneider1991a] V. Sperschneider and G. Antoniou. Logic — A Foundation for Computer

Science. International Computer Science Series. Addison-Wesley, 1991.

[stallings1998a] William Stallings and Richard Van Slyke. Business Data Communications. Pren-

tice Hall, 3 edition, 1998.

[stallings1999a] William Stallings. Cryptography and Network Security: Principles and Practice.

Prentice Hall, 2 edition, 1999.

[stallings2000a] William Stallings. Data and Computer Communications. Prentice Hall, 6 edition,

2000.

[stallings2001a] William Stallings. Business Data Communications. Prentice Hall, 4 edition, 2001. [standefer2001a] Robert Satndefer. Enterprise XML. Morgan Kaufmann Publishers Inc., 2001. [steinmetz1995a] Ralf Steinmetz and Klara Nahrstedt. Multimedia: Computing, Communications

& Applications. Prentice Hall Series in Innovative Technology. Prentice Hall, 1995.

[stevens1990a] W. Richard Stevens. UNIX Network Programming. Prentice Hall Software Series.

Prentice Hall, 1990.

[stevens1992a] W. Richard Stevens. Advanced Programming in the UNIX Environment. Addison-

Wesley Professional Computing Series. Addison-Wesley, 1992.

[stevens1994a] W. Richard Stevens. TCP/IP Illustrated: The Protocols, volume 1 of Addison-

Wesley Professional Computing Series. Addison-Wesley, 1994.

[stevens1996a] W. Richard Stevens. TCP/IP Illustrated: TCP for Transactions, HTTP, NNTP, and the UNIX Domain Protocols, volume 3 of Addison-Wesley Professional Computing Series. Addison-Wesley, 1996.

[stevens1998a] W. Richard Stevens. UNIX Network Programming, Networking APIs: Sockets and

XTI, volume 1. Prentice Hall, 2 edition, 1997.

[stevens1999a] W. Richard Stevens. UNIX Network Programming, Interprocess Communication,

volume 2. Prentice Hall, 2 edition, 1999.

[stone1990a] Harold S. Stone. High-Performance Computer Architecture. Addison-Wesley Series

in Electrical and Computer Engineering. Addison-Wesley, 2 edition, 1990.

[stroustrup1986a] Bjarne Stroustrup. The C++ Programming Language. Addison-Wesley Series

in Computer Science. Addison-Wesley, 1986.

[stroustrup1991a] Bjarne Stroustrup. The C++ Programming Language. Addison-Wesley, 2 edi-

tion, 1991.

[szyperski1998a] Clemens Szyperski. Component Software, Beyond Object-Oriented Programming.

ACM Press/Addison-Wesley, 1998.

[tanenbaum1984a] Andrew S. Tanenbaum. Structured Computer Organization. Prentice Hall, 2

edition, 1984.

[tanenbaum1987a] Andrew S. Tanenbaum. Operating Systems — Design and Implementation.

Prentice Hall, 1987.

[tanenbaum1989a] Andrew S. Tanenbaum. Computer Networks. Prentice Hall, 2 edition, 1989. [tanenbaum1990a] Andrew S. Tanenbaum. Structured Computer Organization. Prentice Hall, 3

edition, 1990.

[tanenbaum1992a] Andrew S. Tanenbaum. Modern Operating Systems. Prentice Hall, 1992. [tanenbaum1996a] Andrew S. Tanenbaum. Computer Networks. Prentice Hall, 3 edition, 1996.

[thai1999a] Thuan L. Thai. Learning DCOM. A Nutshell Handbook. O’Reilly & Associates, Inc.,

1999.

[thai2001a] Thuan Thai and Hoang Q. Lam. .NET Framework Essentials. A Nutshell Handbook.

O’Reilly & Associates, Inc., 2001.

[touretzky1990a] David S. Touretzky. Common Lisp: A Gentle Introduction to Symbolic Compu-

tation. Benjamin/Cummings, 1990.

[tremblett2001a] Paul Tremblett. Instant Enterprise JavaBeans. McGraw-Hill Book Company,

2001.

[tschichold1991a] Jan Tschichold. The Form of the Book. Hartley & Marks, Publishers, 1991. [turner1994a] J. G. Turner and T. L. McCluskey. The Construction of Formal Specifications,

An Introduction to the Model-based and Algebraic Approaches. The McGraw-Hill International Series in Software Engineering. McGraw-Hill Book Company, 1994.

[umar1997a] Amjad Umar. Object-Oriented Client/Server Internet Environments. Prentice Hall,

1997.

[vahalia1996a] Uresh Vahalia. UNIX Internals — The New Frontiers. Prentice Hall, 1996. [valesky1999a] Tom Valesky. Enterprise JavaBeans. Addison-Wesley, 1999. [veen2001a] Jeffrey Veen. The Art & Science of Web Design. New Riders, 2001. [vogel1998a] Andreas Vogel and Keith Duddy. Java Programming with CORBA. John Wiley &

Sons, 2 edition, 1998.

[vogel1999a] Andreas Vogel and Madhavan Rangarao. Programming with Enterprise JavaBeans, JTS and OTS: Building Distributed Transactions with Java and C++. Wiley Computer Publishing. John Wiley & Sons, 1999.

[wall1991a] Larry Wall and Randal L. Schwartz. Programming Perl. A Nutshell Handbook.

O’Reilly & Associates, Inc., 1991.

[wall1996a] Larry Wall, Tom Christiansen, and Randal L. Schwartz. Programming Perl. A Nut-

shell Handbook. O’Reilly & Associates, Inc., 2 edition, 1996.

[walrand1998a] Jean Walrand. Communication Networks. McGraw-Hill Book Company, 2 edition,

1998.

[walsh1994a] Norman Walsh. Making TEX Work. A Nutshell Handbook. O’Reilly & Associates,

Inc., 1994.

[walters1991a] R. F. C. Walters. Categories and Computer Science. Number 28 in Cambridge

Computer Science Texts. Cambridge University Press, 1991.

[wang2001a] Zheng Wang. Internet QoS, Architectures and Mechanisms for Quality of Service.

Morgan Kaufmann Publishers Inc., 2001.

[watters1996a] Aaron Watters, Guido van Rossum, and James C. Ahlstrom. Internet Programming

with Python. M & T Books, 1996.

[wegner2000a] J. D. Wegner. IP Adressing and Subnetting including IPv6. Syngress, 2000. [weiss1998a] Mark Allen Weiss. Data Structures & Problem Solving using Java. Addison-Wesley,

1998.

[weiss1999a] Mark Allen Weiss. Data Structures & Algorithm Analysis in Java. Addison-Wesley, 1999.

[welsh1982a] Jim Welsh and John Elder. Introduction to Pascal. Prentice Hall International Series in Computer Science. Prentice Hall, 2 edition, 1982.

[williams2002a] Hugh E. Williams and David Lane. Web Database Applications with PHP and MySQL. A Nutshell Handbook. O’Reilly & Associates, Inc., 2002.

[wilson1993a] Adrian Wilson. The Design of Books. Chronicle Books, 1993.

[winder2000a] Russel Winder and Graham Roberts. Developing Java Software. ww-cs. John Wiley & Sons, 2 edition, 2000.

[winton2001a] Greg Winton. Palm OS Network Programming. A Nutshell Handbook. O’Reilly & Associates, Inc., 2001.

[wolf1996a] Lars Christian Wolf. Resource Management for Distributed Multimedia Systems. Kluwer Academic Publishers, 1996.

[ziegler1983a] Carol A. Ziegler. Programming System Methodologies. Prentice Hall Software Series. Prentice Hall, 1983.

[zimmermann1996a] Chris Zimmermann, editor. Advances in Object-Oriented Metalevel Architectures and Reflection. CRC Press, 1996.